OpenStudy (mindblast3r):

help pls for teh luv of god! <3

OpenStudy (anonymous):

??

OpenStudy (mindblast3r):

sec i need to write thiis

OpenStudy (anonymous):

ok.

OpenStudy (mindblast3r):

\[\frac{ x(2x+2) }{ x(4x-4)} = \frac{ 2x(x+1) }{ 4x(x-1)}\]

OpenStudy (mindblast3r):

how are these expressions equivalent?

OpenStudy (mindblast3r):

you just factor out the 2 and 4 right?

OpenStudy (anonymous):

ok. all you are doing is like this (a+b)*c= (a*c)+(b*c) like that or you are just moving 2x and 4x out side of the ( )

OpenStudy (mindblast3r):

okay. but can you multiply 2x * 1 = 2 and 4x * 1 = 4??

OpenStudy (anonymous):

there are equivalent because of that

OpenStudy (anonymous):

ik it is the same exprestion just moved out side of the ( )

OpenStudy (anonymous):

that is how they are simmilar.

OpenStudy (mindblast3r):

okay..

OpenStudy (anonymous):

:P lol you dont trust me?

OpenStudy (anonymous):

But you are correct @igreen. @igreen am i correct?

OpenStudy (anonymous):

@answermyquestions

OpenStudy (mindblast3r):

i believe you thank you very much master T :)